OpenStudy (anonymous):

x*x^(10/x) = (x^x)/(x^2) Help please..

OpenStudy (anonymous):

\[x\left(x^{10/x}\right)=\frac{x^x}{x^2}~~?\]

OpenStudy (anonymous):

So i got up to x^2+3x+10=0 but i don't know what I did wrong.

OpenStudy (anonymous):

But this is your equation, right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles \[x\left(x^{10/x}\right)=\frac{x^x}{x^2}~~?\] \(\color{blue}{\text{End of Quote}}\) \[ x^{\frac{10}{x}-1}=x^{x-2}\] And you thought that because the bases are the same, you have \[\frac{10}{x}-1=x-2\] right?

OpenStudy (anonymous):

This is as far as I got:

OpenStudy (anonymous):

The first problem I'm seeing is that you should have \(x^2-3x-10=0\), not plus. The second is that this doesn't give you all the solutions. I'm not quite sure how to get the remaining ones...

OpenStudy (anonymous):

ok thank you.. there's just two more solutions to this

OpenStudy (anonymous):

WolframAlpha says \(\pm1\) are the remaining solutions, but I'm not sure how one would find these without guessing/checking...

OpenStudy (anonymous):

It might involve some logarithms, but that's a wild guess on my part.

OpenStudy (anonymous):

well, thank you for your help.. or else I would stuck thinking about what I did wrong.

OpenStudy (anonymous):

You're welcome!